DSO and Accounts Receivable Harrelson Inc. currently has $765,000 in accounts receivable, and its days sales outstanding (DSO) is 53 days. It wants to reduce its DSO to 35 days by pressuring more of its customers to pay their bills on time. If this policy is adopted, the companys average sales will fall by 20%, what will be the level of accounts receivable following the change? Assume a 365-day year. Round your answer to the nearest cent.
Answer #1

Current Balance in Accounts receivables = $765,000

DSO = 53 days

Total Credit Sales = 765,000*365/53 = $5,268,396 (approx.)

Fall in Sales after new policy = 20%

New Sales = $5,268,396*80% = $4,214,717

Level of Accounts Receivables following the change = $4,214,717*35/365

= $404,150.94